The price of oil is $30 per barrel and the price elasticity is constant and equal to -0.5. An oil embargo reduces the quantity available by 20 percent. Use the arc elasticity formula to calculate the percentage increase in the price of oil.


1
Expert's answer
2022-03-17T09:45:00-0400

Solution.

Initial price "P_0=30"

Price elasticity "E_p=-0.5"

Change in quantity "=20\%=0.2"


New price

"P_1=30\times(1+\frac{0.2}{|-0.5|})"


"P_1=30\times 1.4 =42"


Percentage change in price

"=(\dfrac{42-30}{30})\times 100"


"=40\%"